Calculus I, Fall 2013

Quiz 4

Nov 28, 2013 Show all details.

1. Write down Newton’s method that can be used to find \sqrt[3]{2}. Need not give the numerical value.

2. Solve for y(x) from the equation y′(x) = \frac{1}{1 + 4x^{2}} with y(0) = 1.

3. Evaluate lim_{n→∞} ∑_{k=1}^{n} \frac{1}{n}e^{\frac{k}{n}}.

4. State both parts of Fundamental Theorem of Calculus. Need not prove.

5. Evaluate \frac{d}{dx} ∫_{1}^{x^{2}} e^{t^{2}}dt.

Calculus I, Fall 2013

Quiz 4

Nov 28, 2013 Show all details.

1. Write down Newton’s method that can be used to find \sqrt[3]{2}. Need not give the numerical value.

2. Solve for y(x) from the equation y′(x) = \frac{1}{1 + 4x^{2}} with y(0) = 1.

3. Evaluate lim_{n→∞} ∑_{k=1}^{n} \frac{1}{n}e^{\frac{k}{n}}.

4. State both parts of Fundamental Theorem of Calculus. Need not prove.

5. Evaluate \frac{d}{dx} ∫_{1}^{x^{2}} e^{t^{2}}dt.